# Bourbaki degree of pairs of projective surfaces

### v2

@@ Abstract
The logarithmic tangent sheaf associated to an algebraically independent sequence of homogeneous polynomials - defined as the kernel of the associated Jacobian matrix - naturally generalizes the classical logarithmic tangent sheaf of a divisor in a projective space to the case of subvarieties defined by more than one equation. As is the case for divisors, one may investigate the freeness of such sequences, and other weaker notions.
The present work focuses on sequences of two homogeneous polynomials in four variables. We introduce two positive discrete invariants: the invariant $m$ and the Bourbaki degree of a sequence, inspired by the framework of the Bourbaki degree recently developed for projective plane curves by Jardim-Nejad-Simis. The invariant $m$ plays the role of the Tjurina number of plane projective curves and is bounded by a quadratic relation. We establish results concerning the interplay of minimal degree for syzygies of the Jacobian matrix and the introduced discrete invariants. Our approach uses tools from foliation theory, taking advantage of the fact that the logarithmic sheaf is, up to a twist, the tangent sheaf of a codimension one foliation in $\mathbb{P}^{3}$.
We provide examples and classification results for pencils of cubics and for pairs of a quadric and a cubic polynomials, relating stability and Chern classes with the discrete invariants introduced, while classifying free and nearly-free cases. In particular, one of the nearly-free examples induces an unstable, non-split tangent sheaf for a codimension one foliation of degree $3$, answering, in the negative, a conjecture of Calvo-Andrade, Correa and Jardim from $2018$.
@@ 1 Introduction
The kernel of $\nabla\sigma$, a reflexive sheaf of rank n+1kn+1-k on $\mathbb{P}^{n}$, is called the tangent logarithmic sheaf associated to the sequence $\sigma$. It is considered (see [11]) in analogy with the case of the tangent logarithmic sheaf associated to hypersurfaces in $\mathbb{P}^{n}$. A sequence $\sigma$ is said to be free whenever the sheaf ${\mathcal{T}}_{\sigma}$ splits as a sum of line bundles on $\mathbb{P}^{n}$.
In the case $k=1$ of divisors on $\mathbb{P}^{n}$, one has a short exact sequence of the form:
@@
One should consider sequences rather than the associated subvarieties $X=V(\sigma)$ if $k>1$, because of the following observation: different choices of sequences $\sigma^{\prime},\sigma$ generating the same ideal may have different Jacobian syzygy modules (see [11, Example 2.7]). When $k=1$, however, any different choice $f^{\prime}$ for generator of the ideal $\langle f\rangle$ gives a linear multiple of $\nabla f$, so that ${\mathcal{T}}_{f}\simeq{\mathcal{T}}_{f^{\prime}}$. Similarly, when $d_{1}=\ldots=d_{k}$, elements in sequences $\sigma$, $\sigma^{\prime}$ generating the same ideal vary only by an invertible constant matrix, and in particular ${\mathcal{T}}_{\sigma}\simeq{\mathcal{T}}_{\sigma^{\prime}}$ (see [11, Lemma 2.14]).
We denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ the minimum degree for a non-zero syzygy for the matrix $\nabla\sigma$. Following the terminology of [11], we say that a sequence $\sigma=(f_{1},\ldots,f_{k})$ is compressible if, after a linear change of coordinates, there is a variable that does not occur in any of $f_{1},\ldots,f_{k}$. This is equivalent to $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=0$. Moreover, the number of variables which are independent give trivial copies ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{n}}^{m}\oplus\mathcal{E}$, where $\mathcal{E}$ is a logarithmic sheaf associated to the sequence $\sigma$ in the ring $\kappa[x_{0},\ldots,x_{n-m}]$ (see [11, Lemma 2.8]).
One may also consider the relationship with Bourbaki ideals, as in [17] and [7]. For k=n1k=n-1, a choice of global section of minimum degree $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$ yields a short exact sequence
@@
where $d=-c_{1}({\mathcal{T}}_{\sigma})$ and $B=B_{\nu}\subset\mathbb{P}^{n}$ is a pure codimension two scheme which is generically locally a complete intersection, since ${\mathcal{T}}_{\sigma}$ is a reflexive sheaf of rank two. The scheme $B_{\nu}$ depends on the choice of the syzygy $\nu$, but its Hilbert polynomial is independent of such a choice. In particular, we may consider the degree $\deg(B_{\nu})=\operatorname{Bour}(\sigma)$, which we call Bourbaki degree of the sequence $\sigma$, in analogy with the case studied for $k=1$ and $n=2$ in [17]. By construction, $\operatorname{Bour}(\sigma)=0$ if and only if $\sigma$ is a free sequence. The focus of this work is to understand the above concept for $n=3$ and $k=2$.
@@
and, by definition, the degree $\deg(\Sigma)$ is the leading coefficient above. Moreover, assuming the hypersurface $V(f)$ is reduced, we obtain $s\leq n-2$. For example, if $n=2$, $\deg(\Sigma)$ coincides with the Tjurina number of the projective plane curve $V(f)\subset\mathbb{P}^{2}$, and it appears in the formula of the Bourbaki degree of a projective curve (given in [17]):
@@
Although this is a natural definition for $n=2$, for $n>2$ one has to choose a set of generators for ${\mathcal{T}}_{\sigma}$ to get an ideal sheaf as the cokernel. This problem is addressed in the recent work [7] where the authors consider tame hypersurfaces and other notions of Bourbaki degrees. In this paper, we focus on the case of sequences $\sigma=(f,g)$ on $\mathbb{P}^{3}$, so our logarithmic sheaf ${\mathcal{T}}_{\sigma}$ has also rank two, and we consider the same approach as in [17], since we only have to choose one syzygy to do the construction.
We denote by $\mathcal{Q}_{\sigma}=\operatorname{coker}(\nabla\sigma)$, which corresponds to the coherent sheaf $\mathcal{O}_{J_{f}}(d)$ in the case $k=1$, supported at the singular scheme of $V(f)$. In [11], the authors introduce the Jacobian scheme of $\sigma$, denoted by $\Xi_{\sigma}\doteq V(\bigwedge^{k}\nabla\sigma)$, as the zero locus of the $(k\times k)$-minors of the matrix $\nabla\sigma$. In particular, the reduced support of the sheaf $|\operatorname{supp}\mathcal{Q}_{\sigma}|$ coincides with the reduced support of $\Xi_{\sigma}$, and $c_{1}(\mathcal{Q}_{\sigma})$ coincides with the degree of the greatest common divisor among all $(k\times k)$-minors of $\nabla\sigma$. The generic case is when $c_{1}(\mathcal{Q}_{\sigma})=0$, and this $g.c.d.$ is one, in which case we call $\sigma$ a normal sequence (see 2.1).
If we assume $c_{1}(\mathcal{Q}_{\sigma})=0$, we obtain a Hilbert polynomial of the form
$H(\mathcal{Q}_{\sigma},t)=\frac{\deg(\mathcal{Q}_{\sigma})}{s!}t^{s}+O(t^{s-1}),$
where $s\leq n-2$. We set $m(\sigma)\doteq\deg(\mathcal{Q}_{\sigma})\geq 0$, so $m(\sigma)=0$ if and only if $\Xi_{\sigma}$ has codimension at least three.
In the case $n=3$ and $k=2$, we have sequences $\sigma=(f,g)$ of homogeneous polynomials with degrees $\deg(f)=d_{f}+1,\deg(g)=d_{g}+1$, setting $d\doteq d_{f}+d_{g}$, so our objects are pairs of projective surfaces. We also assume sequences $\sigma$ are normal, as defined above, so that $c_{1}(\mathcal{Q}_{\sigma})=0$ and thus $c_{1}({\mathcal{T}}_{\sigma})=-d$. Since ${\mathcal{T}}_{\sigma}$ has rank two, compressibility implies freeness, and a normal sequence $\sigma$ is compressible if and only if ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d)$.
Using the fact that ${\mathcal{T}}_{\sigma}$ is a rank two reflexive sheaf, we may reproduce analogous considerations as in [17]. We show a formula for the Bourbaki degree of $\sigma$ given by
@@
where $m(\sigma)=\operatorname{ch}_{2}(\mathcal{Q}_{\sigma})$ is the degree of the sheaf $\mathcal{Q}_{\sigma}$ as defined above. The discrete invariant $m(\sigma)$ plays the role of the Tjurina number in the previous context. We compare the different scheme structures between $\operatorname{supp}(\mathcal{Q}_{\sigma})$ and $\Xi_{\sigma}$ in Example 2, considering a free sequence $\sigma=(f,g)$ of polynomials in $\kappa[x_{0},\ldots,x_{3}]$ with $d_{f}=1,d_{g}=2$ and so that $m(\sigma)=5$, but $\deg(\Xi_{\sigma})=6$ and $\deg(\operatorname{supp}(\mathcal{Q}_{\sigma}))=4$ as non-reduced schemes, showing the relationship between $\deg(\mathcal{Q}_{\sigma})$ and $\deg(\operatorname{supp}(\mathcal{Q}_{\sigma}))$ is more complicated than in the case $k=1$, where these two coincide for the sheaf $\mathcal{O}_{J_{f}}(d)$.
About the initial degree and the invariant $m=m(\sigma)$, we show some bounds that are analogous to the case of plane projective curves:
Theorem A. Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials of the ring $\kappa[x_{0},\ldots,x_{3}]$, with degrees $d_{f}+1,d_{g}+1$ respectively. Then:
@@
    $\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq d_{f}+d_{g}$;
@@
We also consider some other inequalities for $m(\sigma)$ and $\operatorname{Bour}(\sigma)$ in Section 3.1 related to freeness and $\mu$-stability of the sheaf ${\mathcal{T}}_{\sigma}$ (see 9): for low enough values of $m(\sigma)$, the sequences are not free, and for high enough values of $\operatorname{Bour}(\sigma)$, the sequences must be $\mu$-stable. We also show the bound $\operatorname{Bour}(\sigma)\leq d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}$, which is attained for example for regular pencils with singular members having only isolated singularities (see Remark 10).
@@
Afterwards, in Section 3.3, we develop relationships between free resolutions for ${\mathcal{T}}_{\sigma}$ and for the ideal sheaf $\mathcal{I}_{B}$, where $B$ is a Bourbaki scheme obtained by the choice of a minimal syzygy. Related to the free resolutions, we introduce the notions of nearly-free sequences and $3$-syzygy sequences. We have the following chain of implications:
$\sigma\text{ is nearly free}\Rightarrow\sigma\text{ is a }3-\text{syzygy sequence}\Rightarrow\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1,$
and the converses do not hold, as we show by building examples with normal pencils of cubics $(d_{f}=d_{g}=2)$.
In Section 4, we explore the structure of codimension one foliation of the sheaf ${\mathcal{T}}_{\sigma}(1)$, presented in [11, Section 9], to obtain characterizations of low initial degrees $\operatorname{indeg}({\mathcal{T}}_{\sigma})\in\{1,2\}$ using the sub-foliations by curves induced by these global sections. For plane projective curves, $e=1$ implies that $\operatorname{Bour}(f)\in\{0,1\}$ (see, for example, [17, Corollary 2.11]). The main theorem of the section is:
Theorem B. Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials with degrees $d_{f}+1,d_{g}+1$, respectively. Then:
@@
In Section 5, we finish with a study on two particular families of normal sequences: pencils of cubics $(d_{f}=d_{g}=2)$ and sequences with $d_{f}=1,d_{g}=2$, defining a degree $6$ curve inside a quadric surface in $\mathbb{P}^{3}$. In these two classes, we classify all free and nearly free cases in terms of their discrete invariants and establish some stability results, which are summarized below:
Theorem C. Let $\sigma=(f,g)$ be a normal pencil of cubic surfaces in $\mathbb{P}^{3}$. Then, if we denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$:
@@
In this case, we have an example of a strictly $\mu$-semistable logarithmic sheaf ${\mathcal{T}}_{\sigma}$ with $m(\sigma)=4$ (see Example 16), but the bound above for stability may not be sharp.
Theorem D. Let $\sigma=(f,g)$ be a normal sequence with $d_{f}=1,d_{g}=2$. Then, if $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$:
@@
    If $m(\sigma)\leq 3$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
Example 15 induces a codimension one foliation $\mathcal{F}$ of degree $3$ with tangent sheaf $T_{\mathcal{F}}={\mathcal{T}}_{\sigma}(1)$ which is non-split and not $\mu$-semistable. This example provides a negative answer to a conjecture posed by Calvo-Andrade, Correa and Jardim, in [2]:
@@
Along the text, we describe examples developed computationally with aid of Macaulay2 software ([12]) for the two families of sequences.

### v3

@@ Abstract
The logarithmic tangent sheaf associated to an algebraically independent sequence of homogeneous polynomials — defined as the kernel of the associated Jacobian matrix — naturally generalizes the classical logarithmic tangent sheaf of a divisor in a projective space to the case of subvarieties defined by more than one equation. As is the case for divisors, one may investigate the freeness of such sequences, and other weaker notions.
The present work focuses on sequences of two homogeneous polynomials in four variables. We introduce two positive discrete invariants: the invariant $m$ and the Bourbaki degree of a sequence, inspired by the framework of the Bourbaki degree recently developed for projective plane curves by Jardim-Nejad-Simis. The invariant $m$ plays the role of the Tjurina number of plane projective curves and is bounded by a quadratic relation of the degrees. We establish results concerning the interplay of minimal degree for syzygies of the Jacobian matrix and the introduced discrete invariants. Our approach uses tools from foliation theory, taking advantage of the fact that the logarithmic sheaf is, up to a twist, the tangent sheaf of a codimension one foliation in $\mathbb{P}^{3}$.
We provide examples and classification results for pencils of cubics and for pairs of a quadric and a cubic polynomials, relating stability and Chern classes with the discrete invariants introduced, while classifying free and nearly free cases. In particular, one of the nearly free examples induces an unstable, non-split tangent sheaf for a codimension one foliation of degree $3$, answering, in the negative, a conjecture of Calvo-Andrade, Correa and Jardim from $2018$.
@@ 1 Introduction
The kernel of $\nabla\sigma$, a reflexive sheaf of rank n+1kn+1-k on $\mathbb{P}^{n}$, is called the tangent logarithmic sheaf associated to the sequence $\sigma$. This definition is in analogy (see [undefj]) with the case of the tangent logarithmic sheaf associated to divisors in $\mathbb{P}^{n}$, namely for $k=1$. A sequence $\sigma$ is said to be free whenever the sheaf ${\mathcal{T}}_{\sigma}$ splits as a sum of line bundles on $\mathbb{P}^{n}$.
In the case of divisors on $\mathbb{P}^{n}$, where $k=1$, one has a short exact sequence of the form:
@@
For the generalization of $k>1$, one should consider sequences rather than the associated ideals $I=(f_{1},\ldots,f_{k})$, as pointed out by [undefj, Example 2.7]. When $d_{1}=\ldots=d_{k}$, elements in sequences $\sigma$, $\sigma^{\prime}$ generating the same ideal vary only by an invertible constant matrix, and furthermore ${\mathcal{T}}_{\sigma}\simeq{\mathcal{T}}_{\sigma^{\prime}}$ (see [undefj, Lemma 2.14]).
We denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ the minimum degree for a non-zero syzygy for the matrix $\nabla\sigma$, called the initial degree. Following the terminology of [undefj], we call a sequence $\sigma=(f_{1},\ldots,f_{k})$ compressible if, after a linear change of coordinates, there is a variable that does not occur in any of $f_{1},\ldots,f_{k}$. This is equivalent to $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=0$. Moreover, the number of variables which are independent give trivial copies ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{n}}^{m}\oplus\mathcal{E}$, where $\mathcal{E}$ is a logarithmic sheaf associated to the sequence $\sigma$ in the ring $\kappa[x_{0},\ldots,x_{n-m}]$ (see [undefj, Lemma 2.8]).
One may also consider the relationship with Bourbaki ideals, as in [undefr] and [undeff]. For k=n1k=n-1, a choice of global section of minimum degree $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$ yields a short exact sequence
@@
where $B=B_{\nu}\subset\mathbb{P}^{n}$ is a pure codimension two scheme which is generically locally a complete intersection. The scheme $B_{\nu}$ depends on the choice of the syzygy $\nu$, but its Hilbert polynomial is independent of such a choice. In particular, we may consider the degree $\deg(B_{\nu})=\operatorname{Bour}(\sigma)$, which we call Bourbaki degree of the sequence $\sigma$, in analogy with the case studied for $k=1$ and $n=2$ in [undefr]. By construction, $\operatorname{Bour}(\sigma)=0$ if and only if $\sigma$ is a free sequence. The focus of this work is to understand the concept above for $n=3$ and $k=2$.
@@
and, by definition, the degree $\deg(\Sigma)$ is the leading coefficient above. Moreover, assuming the hypersurface $V(f)$ is reduced, we obtain $s\leq n-2$. For example, if $n=2$, $\deg(\Sigma)$ coincides with the Tjurina number of the projective plane curve $V(f)\subset\mathbb{P}^{2}$, and it appears in the formula of the Bourbaki degree of a projective curve (given in [undefr]):
@@
Although this is a natural definition for $n=2$, for $n>2$ one has to choose a set of generators for the logarithmic sheaf ${\mathcal{T}}_{f}$ to get an ideal sheaf as the cokernel. This problem is addressed in a recent work [undeff] where the authors consider another kind of Bourbaki sequences, recently developed in [undefq], by choosing generators of maximal degree.
We denote by $\mathcal{Q}_{\sigma}=\operatorname{coker}(\nabla\sigma)$ the cokernel of the Jacobian matrix, which corresponds to the coherent sheaf $\mathcal{O}_{J_{f}}(d)$ in the case $k=1$, supported at the singular scheme of $V(f)$. In [undefj], the authors introduce the Jacobian scheme of $\sigma$, denoted by $\Xi_{\sigma}\doteq V(\bigwedge^{k}\nabla\sigma)$, as the zero locus of the $(k\times k)$-minors of the matrix $\nabla\sigma$. In particular, the reduced support of the sheaf $|\operatorname{supp}\mathcal{Q}_{\sigma}|$ coincides with the reduced support of $\Xi_{\sigma}$. The generic case is when $\dim(\mathcal{Q}_{\sigma})\leq n-2$, and in particular $c_{1}(\mathcal{Q}_{\sigma})=0$. In this case, we call $\sigma$ a normal sequence.
For now on, we assume $n=3$ and $k=2$, that is, sequences $\sigma=(f,g)$ of homogeneous polynomials $f,g\in R\doteq\kappa[x_{0},\ldots,x_{3}]$ with degrees $\deg(f)=d_{f}+1,\deg(g)=d_{g}+1$, and we set $d\doteq d_{f}+d_{g}$. We also assume sequences $\sigma$ are normal, as defined above, so that $c_{1}(\mathcal{Q}_{\sigma})=0$ and $c_{1}({\mathcal{T}}_{\sigma})=-d$. Since the logarithmic sheaf ${\mathcal{T}}_{\sigma}$ has rank two, compressibility implies freeness, and a normal sequence $\sigma$ is compressible if and only if ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d)$.
If $\sigma$ is a normal sequence, $\dim(\mathcal{Q}_{\sigma})\leq 1$ and the Hilbert polynomial is of the form
$H(\mathcal{Q}_{\sigma},t)=\deg(\mathcal{Q}_{\sigma})t+O(1),$
We set $m(\sigma)\doteq\deg(\mathcal{Q}_{\sigma})\geq 0$, so $m(\sigma)=0$ if and only if $\Xi_{\sigma}$ is zero-dimensional. From the associativity formula applied to the graded homogeneous module $\mathcal{Q}$ associated to $\mathcal{Q}_{\sigma}$, we obtain the expression (see, for example, [undefs, Theorem $14.7$])
$m(\sigma)=\sum_{P\in\operatorname{Ann}(\mathcal{Q}_{\sigma}),\dim(R/P)=2}\operatorname{{length}}_{R_{P}}\mathcal{Q}_{P}\cdot\deg(R/P).$
Using the fact that ${\mathcal{T}}_{\sigma}$ is a rank two reflexive sheaf, we reproduce analogous considerations as in [undefr]. We show a formula for the Bourbaki degree of $\sigma$ given by
@@
so that $\operatorname{Bour}(\sigma)=0$ if and only if $\sigma$ is a free sequence. A first result of the paper is some bounds for the discrete quantities defined above.
Theorem A(Theorem A) Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials of the ring $R$, with degrees $d_{f}+1,d_{g}+1$ respectively, and set $d=d_{f}+d_{g}$. Then:
@@
    $\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq d$;
@@
We also consider some other inequalities for $m(\sigma)$ and $\operatorname{Bour}(\sigma)$ in Section 2.1 related to freeness and $\mu$-stability of the sheaf ${\mathcal{T}}_{\sigma}$ (see 8): for low enough values of $m(\sigma)$, the sequences are not free, and for high enough values of $\operatorname{Bour}(\sigma)$, the sequences must be $\mu$-stable. In general, the Bourbaki degree is bounded by $\operatorname{Bour}(\sigma)\leq d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}$, and the maximum is attained precisely by the matrices $\nabla\sigma$ for which the Buchsbaum–Rim complex is a minimal resolution, or equivalently $m(\sigma)=0$. Geometrically, all regular pencils ($d_{f}=d_{g}$) with isolated singularities satisfy this property (see Remark 9).
@@
Afterwards, in Section 2.3, we develop relationships between free resolutions for ${\mathcal{T}}_{\sigma}$ and for the ideal sheaf $\mathcal{I}_{B}$, where $B$ is the Bourbaki scheme obtained by the choice of a minimal syzygy. Related to the free resolutions, we introduce the notions of nearly free sequences and $3$-syzygy sequences, with the following chain of implications
$\sigma\text{ is nearly free}\Rightarrow\sigma\text{ is a }3\text{-syzygy sequence}\Rightarrow\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1.$
The converses do not hold, as we show by building examples with normal pencils of cubics $(d_{f}=d_{g}=2)$. We point out nearly free sequences are never locally free, as we show in 19.
In Section 3, we explore the structure of codimension one foliation of the sheaf ${\mathcal{T}}_{\sigma}(1)$, presented in [undefj, Section 9], to obtain characterizations of low initial degrees $\operatorname{indeg}({\mathcal{T}}_{\sigma})\in\{1,2\}$ using the sub-foliations by curves induced by these global sections. For plane projective curves, $e=1$ implies that $\operatorname{Bour}(f)\in\{0,1\}$ (see, for example, [undefr, Corollary 2.11]). The main theorem of the section is:
Theorem B.(Theorem B) Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials with degrees $d_{f}+1,d_{g}+1$, respectively. Then:
@@
In Section 4, we finish with a study on two particular families of normal sequences: pencils of cubics $(d_{f}=d_{g}=2)$ and sequences with $d_{f}=1,d_{g}=2$, defining a degree $6$ curve inside a quadric surface in $\mathbb{P}^{3}$. To do so, we use classical results on bounds for Chern classes of reflexive sheaves in $\mathbb{P}^{3}$. We classify all free and nearly free cases in these classes in terms of their discrete invariants and establish some stability results, which are summarized below:
Theorem C.(Theorem C)Let $\sigma=(f,g)$ be a normal pencil of cubic surfaces in $\mathbb{P}^{3}$. Then, if we denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$:
@@
In this case, we have an example of a strictly $\mu$-semistable logarithmic sheaf ${\mathcal{T}}_{\sigma}$ with $m(\sigma)=4$ (see Example 15), but the bound given above for stability may not be sharp.
Theorem D.(Theorem D) Let $\sigma=(f,g)$ be a normal sequence with $d_{f}=1,d_{g}=2$. Then, if $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$:
@@
    If $m(\sigma)=3$ then $\operatorname{Bour}(\sigma)=2$, with a possible unstable case $(c_{3}=6)$ and stable cases with $c_{3}\in\{0,2,4\}$. Among these, we have examples for the stable cases with $c_{3}=2$ (see Example 58) and $c_{3}=4$ (see Example 59).
  • (e)
    If $m(\sigma)<3$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
Example 14 induces a codimension one foliation $\mathcal{F}$ of degree $3$ with tangent sheaf $T_{\mathcal{F}}={\mathcal{T}}_{\sigma}(1)$ which is non-split and not $\mu$-semistable. This example provides a negative answer to a conjecture posed by Calvo-Andrade, Correa and Jardim, in [undefa]:
@@
Along the text, we describe examples developed computationally with aid of Macaulay2 software ([undefl]) for the two families of sequences.